# ARM: Refining Multivariate Forecasting with Adaptive Temporal-Contextual Learning

### v1

@@ Abstract
Long-term time series forecasting (LTSF) is important for various domains but is confronted by challenges in handling the complex temporal-contextual relationships. As multivariate input models underperforming some recent univariate counterparts, we posit that the issue lies in the inefficiency of existing multivariate LTSF Transformers to model series-wise relationships: the characteristic differences between series are often captured incorrectly. To address this, we introduce ARM: a multivariate temporal-contextual adaptive learning method, which is an enhanced architecture specifically designed for multivariate LTSF modelling. ARM employs Adaptive Univariate Effect Learning (AUEL), Random Dropping (RD) training strategy, and Multi-kernel Local Smoothing (MKLS), to better handle individual series temporal patterns and correctly learn inter-series dependencies. ARM demonstrates superior performance on multiple benchmarks without significantly increasing computational costs compared to vanilla Transformer, thereby advancing the state-of-the-art in LTSF. ARM is also generally applicable to other LTSF architecture beyond vanilla Transformer.
@@ 1 Introduction
Long-term time series forecasting (LTSF) is a critical task across various fields such as finance, epidemiology, electricity, and traffic, aiming to predict future values over an extended horizon, thereby facilitating optimal decision-making, resource allocation, and strategic planning (Martínez-Álvarez et al., 2015; Lana et al., 2018; Kim, 2003; Yang et al., 2015; Ma et al., 2022). However, modeling LTSF poses numerous challenges due to the complex and entangled characteristics of multivariate time series data, often leading to overfitting and erroneous pattern learning, thereby compromising model performance (Peng & Nagata, 2020; Cao & Tay, 2003; Sorjamaa et al., 2007).
Recently, Transformers (Vaswani et al., 2017) have significantly outperformed other structures in sequential modeling. They have showcased advancements in LTSF modeling (Zhou et al., 2022; Wu et al., 2022; Zhou et al., 2021; Li et al., 2019). While models with multivariate time series inputs are generally considered effective for capturing both temporal and contextual relationship, recent studies have shown that LTSF models with univariate input can surprisingly outperform their multivariate counterparts (Zeng et al., 2022; Nie et al., 2023). This is counter-intuitive, as univariate models are limited in capturing relationships across multiple series, which is crucial for LTSF.
In this study, we argue that existing multivariate LTSF Transformers fall short in properly modeling series-wise relationships due to suboptimal training and data processing methods, leading to subpar performance compared to univariate models. These models struggle to handle significant differences of characteristic across various input series, such as the differences in temporal dependencies, differences in local temporal patterns, and differences in series-wise dependencies beyond intra-series relationships. Improper mixing of these distinct series within crucial components of the Transformer blocks, such as temporal attention and input embedding, undermines the model’s capacity to differentiate them effectively during forecasting tasks.
@@
Figure 1: Three intuitive problems arising from wrongly handling input series with characteristic differences. (a) Adaptive Univariate Effect Learning of Output Mean (see §3.1): We illustrate how different methods estimate the output mean for five example time series. The blue lines denote the original series, with the input and output sections separated by a purple line. Using yellow and red lines, we represent methods that employ the input mean (RevIN) and the last input value (NLinear) respectively as the output mean. These previous methods are suboptimal as they don’t account for varying lookback lengths required for different series. We highlight the ideal lookback areas with green boxes. The Adaptive EMA in our method, represented by a green line, effectively learns the optimal output mean, thereby handling series with diverse temporal dependencies more accurately. (b) The Necessity of Building Reasonable Multivariate Temporal Representation (see §3.3), which is demonstrated by five series from the Electricity dataset. A black vertical line is used to indicate the current time step $t$. Their associated temporal relationships and local patterns vary for different series. Thus, adopting the values from $X^{(t)}$ as the representation for time step $t$, as done in previous models, wrongly blends the modeling across series with distinct characteristics. We use green, red, and yellow boxes to depict the optimal scope of local patterns that should be considered for different series at time $t$. To accommodate these distinct patterns, we introduce a multi-window local convolutional module to construct reasonable representation for multivariate inputs. (c) The Inability of Previous Models to Learn Obvious Inter-Series Dependencies: We illustrate the ineffectiveness of existing multivariate LTSF Transformers using a dataset ”Multi”, where the subsequent three series are simple shifts of the first (details in section A.2). A ”copy-paste” operation should suffice for forecasting, yet previous multivariate models, such as Autoformer, fall short to accomplish this task efficiently. In Figure 5, we visually demonstrate ARM’s enhancement in handling these strong inter-series relationships for LTSF models.
@@ 2 Related Works
Time series forecasting has long been a popular research topic. Traditional approaches, such as the ARIMA model (Box et al., 1974) and the Holt-Winters seasonal method (Holt, 2004), provided theoretical guarantees but were limited for complex time series data. Deep learning models (Oreshkin et al., 2019; Sen et al., 2019) introduced a new era in this field. RNNs (Hochreiter & Schmidhuber, 1997; Wen et al., 2017; Rangapuram et al., 2018; Shih et al., 2019; Salinas et al., 2020; Qin et al., 2017) allowed for the summary of past information in compact internal memory states, updated recursively with new inputs at each time step. CNNs and temporal convolutional networks (TCNs) (Lai et al., 2018; Borovykh et al., 2017; van den Oord et al., 2016) further advanced the field, capturing local temporal features effectively, though with limitations on long-term dependencies. For LTSF, we have summarized the methods of data processing used by previous models in Figure 2.
Recently, Transformer (Vaswani et al., 2017) succeed in sequential modeling tasks with the effective attention mechanism. Transformer derivatives like LogTrans (Li et al., 2019) introduced local convolution, reducing complexity for LTSF. Similarly, Informer (Zhou et al., 2021) and Autoformer (Wu et al., 2022) extended Transformer with efficient attention and auto-correlation mechanisms respectively, while FedFormer (Zhou et al., 2022) and PyraFormer (Liu et al., 2021) utilized improved attention to achieve lower complexity. However, a recent study (Zeng et al., 2022) argued that Transformer may fail to correctly understand multivariate time series structure. PatchTST (Nie et al., 2023) address the issue with independent input but fell short in modeling multivariate correlations. CrossFormer (Zhang & Yan, 2023) tried to address with 2D attention, but its hierarchical segmentation failed adapt length per series. To address the challenges, we propose ARM to enhance the accuracy and adaptability of LTSF with proper training and data processing methods.
Figure 2: Different types of LTSF models. Here, we present six types of LTSF modelling architectures, categorized into univariate and multivariate forecasting methods. The parallel arrows represent the independent processing of each series, while converging arrows indicate the mixing of different series together in the model. The univariate methods, labeled as (a), (b), and (c), process each series independently, without consideration for inter-series relationships. Method (a) trains separate models for each series, represented by models such as individual DLinear; Method (b) uses a single, shared-parameter model for all series, exemplified by models like PatchTST; Method (c) dynamically selects best models for individual input series as outlined in §3.1. The multivariate methods, designated as (d), (e), and (f), aim to capture inter-series dependencies. Method (d) employs a standard multivariate structure but falls short in disentangling inter-series relationships, hence being outperformed by univariate methods. Preexisting models include Informer, Autoformer, etc.; Method (e) augments univariate models with multivariate factors, adding computational complexity; Method (f) features our proposed Random Dropping training strategies (see §3.2), enabling effective learning of inter-series dependencies without extra computational cost.
@@ 3.1 Module A: Adaptive Univariate Effect Learning
In previous LTSF Transformers such as Informer and Autoformer (Wu et al., 2022; Zhou et al., 2021), the $L_{P}$ region of the decoder input is filled with 0 values as default outputs (models without an $L_{P}$ part input can also be considered to use 0 default). However, the actual meaning of 0 can vary significantly across series due to their distributional differences, making it difficult for multivariate models to determine the output level for each series. As the output level predominantly influences the loss like MSE during training, the model tends to use the majority of parameters to learn the output level, making it challenging to capture finer-grained temporal patterns and inter-series dependencies.
@@
Recent studies tried to mitigate the aforementioned issues by focusing on “eliminating distribution shifts between input and output”. These approaches attempt to estimate the mean and variance, eliminating them from the model input, and restoring these values at the output. RevIN (Kim et al., 2021) calculates the mean and variance for each input series while NLinear (Zeng et al., 2022) simply employs the last value of each series at $L_{I}$ as the output mean. However, as illustrated in Figure 1 (a), for series with differing statistical properties, the lookback lengths required to determine their output distribution tend not to be constant. Relying solely on last or all input values will weaken performance. Thus, we propose an adaptive method based on a learnable exponential moving average (EMA), which dynamically adjusts the output mean and variance for different series.
@@
Previous methods still employ constant value to fill the $L_{P}$ part of each series without building temporal patterns. In multivariate LTSF, handling the univariate temporal patterns prior to the multivariate modeling can help to learn inter-series relationships more accurately. In multivariate datasets, suppose a given series $i$ has no causal relationships with other series. If we utilize a univariate model to initialize $X_{P}^{i}$ part based on $X_{I}^{i}$ before other operations, this preliminary step will allow the model to focus more on capturing other inter-series relationships. Previous works like LSTNet (Lai et al., 2018) and ES-RNN (Smyl, 2020) have used classic autoregressive and exponential smoothing models for this purpose. However, both methods are limited in capturing basic patterns and suffer from poor generalization. Therefore, we use an adaptive method based on the Mixture of Experts (MoE) for temporal patterns to better handle these complexities, as illustrated in Figure 2 (c).
Figure 3: Overall Architecture of ARM with Vanilla Transformer as encoder-decoder. The left side depicts the global workflow incorporating the AUEL, Random Dropping, and MKLS modules. The right side illustrates the specific computational process employed by the model when training on a given multivariate time series input.
@@
where $\circ$ denotes element-wise multiplication, $\bm{t}=[1,2,\cdots,L_{I}]$ is a vector representing the timesteps, and $[\cdot]_{j}$ indicates the $j$-th element of the vector enclosed in brackets; $X_{I}^{(L_{I}-w_{j}:L_{I}),i}$ is a vector composed of the part of input series $X_{I}^{i}$ that goes back $w_{j}$ steps from the last value. When the $\alpha^{i}$ of channel $i$ approaches 1, $p_{global}^{i}=1$, and $p_{local}^{i}=0$, the AUEL of output distribution becomes RevIN. When $\alpha^{i}$ approaches 0, the AUEL will be similar to NLinear which uses last values.
AUEL of Temporal Patterns   Figure 2 parts (a) and (b) illustrate two types of univariate input LTSF models: (b) with fully-shared parameters fits datasets with similar series characteristics while (a) with independent channels is suited data with diverse series. Typically, a dataset may contain multiple clusters of similar series. Thus, (c), an intermediary approach between (a) and (b) is preferable. We use MoE, simillar to the module in Switch Transformer (Fedus et al., 2022), to dynamically select the predictor based on series characteristics to initialize the temporal patterns for each series (see A.5.1 for more details on the implementation of MoE for temporal pattern learning).
[…]
$\widetilde{X}^{*i}=\gamma^{i}\left(\frac{X^{i}-E^{i}}{S^{i}+\epsilon}\right)+\beta^{i},\ \ \ \ \widetilde{X}^{i}=\left[\widetilde{X}^{*i}_{I}\ \ \ \ \mathtt{MoE}\left(\left[\widetilde{X}^{*i}_{I}\ \ \ \ \widetilde{X}^{*i}_{P}\right]\right)\right]\vspace{-5pt}$ (2)
@@ 3.2 Module R: Random Dropping and Inter-Series Dependency Learning
In LTSF, accurately modeling inter-series dependencies is often challenging due to the simultaneous handling of complex temporal and contextual information. Figure 1 (c) provides an illustrative example where a multivariate Transformer fails to capture some very evident causal dependencies.
@@
We introduce a Random Dropping strategy to train multivariate LTSF models effectively. This strategy facilitates the learning of inter-series dependencies by efficiently decoupling relationships between different series, without increasing computational complexity. Applied during the training phase, Random Dropping simultaneously sets a random subset of series to zero in both the input and training target, as illustrated in Figure 2 (f). The model thus learns the contributions of the current subset to the forecasting of series within that subset. Through this random selection, the model incrementally discerns which series most significantly influence the forecasting of others, thereby effectively mitigating the risk of overfitting. Conceptually, Random Dropping can be viewed as a form of model ensemble that constructs a pool of forecasting models for every possible subset-to-subset relationship, then enabling the more effective models among them. Utilizing Random Dropping can straightforwardly enhance the learning ability of multivariate LTSF models in terms of inter-series dependency, achieving disentanglement between series.
@@ 3.3 Module M: Multi-kernel Local Smoothing
In order to enhance the understanding of multivariate temporal structure, We propse the Multi-kernel Local Smoothing (MKLS) block which is used in conjunction with the Transformer blocks, as shown in Figure 4. MKLS uses multiple different 1D convolutional kernels and a channel-wise attention to learn and extract local information with adjustable view length. To address the aforementioned challenges of building reasonable temporal representations and enhancing locality, we provide two usage of MKLS. (i) Pre-MKLS is applied before feeding data into the Transformer blocks, helping build meaningful local representations for multivariate input. (ii) Post-MKLS processes data in parallel with the Transformer blocks, playing a role like local attention with adjustable local windows. Through the incorporation of channel-wise attention, MKLS learns different kernel weights, providing robust adaptability for series with strong characteristic differences.
@@
MKLS block   Let the input for the MKLS block be $X\in\mathbb{R}^{l\times d}$, where $l$ is the length of the temporal dimension and $d$ is the number of channels. Suppose we set $n$ 1D convolutional layers with kernel sizes $s_{j}\in\left(s_{1},\cdots,s_{n}\right)$. First, we obtain the computation results $\widetilde{X}_{j}$ from the $j$-th convolutional layer by: $\widetilde{X}_{j}=\mathtt{Conv1d}_{j}\left(X_{j}\right)$. To acquire the weights $m^{i}_{j}$ for kernel $j$ on channel $i$, we first perform kernel dropout and average-pooling on $\widetilde{X}_{j}$, and then calculate the cross attention between the averaged output and the original $X$. For kernel dropout, we randomly drop a certain proportion $r_{k}$ of kernel outputs, which enable independent learning of local patterns thus prevent overfitting. Let the tensor $\widetilde{\mathcal{X}}\in\mathbb{R}^{l\times d\times n}$ contain the results of each $\widetilde{X}_{k}\in\mathbb{R}^{l\times d}$. We now perform average-pooling to obtain another input $\widetilde{X}_{\mathtt{Avg}}=\frac{1}{1-r_{k}}\cdot\frac{\sum_{k=1}^{n}\widetilde{X}_{k}}{n}$ for the attention calculation besides original $X$. Then, we use $\widetilde{X}_{\mathtt{Avg}}$ and $X$ as Q and K in the attention, and obtain the kernel weights $M\in\mathbb{R}^{d\times n}$ for each channel using a trainable latent V. The specific calculation is: $M=\mathtt{softmax}\left(\frac{W_{Q}(X^{\top}+P_{E})\left(W_{K}(\widetilde{X}^{\top}_{\mathtt{Avg}}+P_{E})\right)^{\top}}{\sqrt{d}}\right)W_{V}$ , where $W_{Q}\in\mathbb{R}^{d\times d}$ and $W_{K}\in\mathbb{R}^{d\times d}$ are the weights for the Q and K in the attention, initialized as identity matrices to retain the original channel structure before training; $W_{V}\in\mathbb{R}^{d\times n}$ is designed to be a latent matrix, which projects the cross-channel attention map to the dimension of $n$, obtaining the kernel weights $M$; $P_{E}$ is the trainable positional embedding matrix. Now, we can calculate the weighted kernel output $\widetilde{S}$ as: $\widetilde{S}_{t,i}=\sum_{j=1}^{n}\left(\widetilde{\mathcal{X}}_{t,i,j}*M_{i,j}\right)_{t,i,j}+X_{t,i}$, where *** represents the element-wise multiplication performed on the channel and multi-kernel dimensions; the output $\widetilde{S}_{t,i}$ is the local smoothed output after weighting $n$ local kernels. A residual shortcut of the input is added and denoted by $X_{t,i}$.
Figure 4: Structure of Multi-kernel Local Smoothing (MKLS). The central part of the figure illustrates the detailed computation of MKLS, incorporating multiple 1D convolutions and channel-wise attention. The left side and right side presents the application method of Pre-MKLS and Post-MKLS, respectively. Kernel sizes here are $s_{1},s_{2},s_{3}=(7,24,96)$
@@
Pre-MKLS and Post-MKLS   MKLS module, used alongside Transformer blocks, effectively manages local shape extraction and alignment in multivariate input. Pre-MKLS constructs local smoothed representations, offering more self-adjustment based on input compared to the pure 1D convolutional approach like (Li et al., 2019). Post-MKLS, processing data in parallel with the Transformer, enables the latter to focus more on learning long-term connections, with its logic similar to the local attention methods such as (Child et al., 2019; Roy* et al., 2020) and parallel local convolutional layer in (Wu et al., 2020). In comparison to these local attention methods, Post-MKLS can dynamically adjust local window weight for different input data to modulate the local view. Pre-MKLS is defined as: $\mathtt{PreMKLS}(X,K,V):=\mathtt{Cross}(Q,K,V)$, where $Q=\mathtt{MKLS}(X)+X+P_{E}$; MKLS represents the computation described in the previous section; $P_{E}$ is the trainable positional embedding; and the Cross module can be regarded as Transformer Encoders with replaceable self-attention or cross-attention. If using self-attention, then $Q=K=V$, otherwise, $K$ and $V$ are externally provided matrices. If the last dimension lengths of $Q$, $K$, and $V$ are different, then we adjust the $W$ projection matrices of $Q$, $K$, and $V$ to provide a aligned dimension. Post-MKLS is defined as: $\mathtt{PostMKLS}(X,K,V):=\mathtt{LayerNorm}(\mathtt{MKLS}(X)+\mathtt{Cross}(Q,K,V))$, where $Q=X+P_{E}$.

### v2

@@ Abstract
Long-term time series forecasting (LTSF) is important for various domains but is confronted by challenges in handling the complex temporal-contextual relationships. As multivariate input models underperforming some recent univariate counterparts, we posit that the issue lies in the inefficiency of existing multivariate LTSF Transformers to model series-wise relationships: the characteristic differences between series are often captured incorrectly. To address this, we introduce ARM: a multivariate temporal-contextual adaptive learning method, which is an enhanced architecture specifically designed for multivariate LTSF modelling. ARM employs Adaptive Univariate Effect Learning (AUEL), Random Dropping (RD) training strategy, and Multi-kernel Local Smoothing (MKLS), to better handle individual series temporal patterns and effectively learn inter-series dependencies. ARM demonstrates superior performance on multiple benchmarks without significantly increasing computational costs compared to vanilla Transformer, thereby advancing the state-of-the-art in LTSF. ARM is also generally applicable to other LTSF architecture beyond vanilla Transformer. The code implementation is available at this link.
@@ 1 Introduction
Long-term time series forecasting (LTSF) is a critical task across various fields such as finance, epidemiology, electricity, and traffic, aiming to predict future values over an extended horizon, thereby facilitating optimal decision-making, resource allocation, and strategic planning (Martínez-Álvarez et al., 2015; Lana et al., 2018; Kim, 2003; Yang et al., 2015; Ma et al., 2022). However, modeling LTSF poses numerous challenges due to the complex and entangled characteristics of multivariate time series data, often leading to overfitting and erroneous pattern learning, thereby compromising model performance (Peng and Nagata, 2020; Cao and Tay, 2003; Sorjamaa et al., 2007).
Recently, Transformers (Vaswani et al., 2017) have significantly outperformed other structures in sequential modeling. They have showcased advancements in LTSF modeling (Zhou et al., 2022; Wu et al., 2021; Zhou et al., 2021; Li et al., 2019). While models with multivariate time series inputs are generally considered effective for capturing both temporal and contextual relationship, recent studies have shown that LTSF models with univariate input can surprisingly outperform their multivariate counterparts (Zeng et al., 2023; Nie et al., 2022). This is counter-intuitive, as univariate models are limited in capturing relationships across multiple series, which is crucial for LTSF.
In this study, we argue that existing multivariate LTSF Transformers fall short in properly modeling series-wise relationships due to suboptimal training and data processing methods, leading to subpar performance compared to univariate models. These models struggle to handle significant differences of characteristic across various input series, such as the differences in temporal dependencies, differences in local temporal patterns, and differences in series-wise dependencies beyond intra-series relationships, as exemplified in Figure 1. Improper mixing of these distinct series within crucial components of the Transformer blocks, such as temporal attention and input embedding, undermines the model’s capacity to differentiate them effectively during forecasting tasks.
@@
Figure 1: Three intuitive problems arising from wrongly handling input series with characteristic differences (further explained in §A.3). (a) Adaptive estimation of output mean necessity for series with diverse characteristics (see §3.1). The Adaptive EMA (green) in our method outperforms previous approaches like RevIN (yellow) and NLinear (red) by adapting to varied temporal dependencies. (b) The necessity of building reasonable multivariate temporal representation (see §3.3). Our multi-window local convolutional module addresses the inadequacy of former models in modeling different local patterns across series. (c) The Inability of Previous Models to Learn Obvious Inter-Series Dependencies: Demonstrated using the ”Multi” dataset generated with simple shifting (see A.2), existing models like Autoformer fail to learn this simple ”copy-paste” operations. See Figure 6 for the visualization to show our ARM’s performance boost in this dataset.
@@ 2 Related Works
Time series forecasting has long been a popular research topic. Traditional approaches, such as the ARIMA model (Box et al., 1974) and the Holt-Winters seasonal method (Holt, 2004), provided theoretical guarantees but were limited for complex time series data. Deep learning models (Oreshkin et al., 2020; Sen et al., 2019) introduced a new era in this field. RNNs (Hochreiter and Schmidhuber, 1997; Wen et al., 2017; Rangapuram et al., 2018; Shih et al., 2019; Salinas et al., 2020; Qin et al., 2017) allowed for the summary of past information in compact internal memory states, updated recursively with new inputs at each time step. CNNs and temporal convolutional networks (TCNs) (Lai et al., 2018; Borovykh et al., 2017; van den Oord et al., 2016) further advanced the field, capturing local temporal features effectively, though with limitations on long-term dependencies.
Recently, Transformer (Vaswani et al., 2017) succeed in sequential modeling tasks with the effective attention mechanism. Transformer derivatives like LogTrans (Li et al., 2019) introduced local convolution, reducing complexity for LTSF. Similarly, Informer (Zhou et al., 2021) and Autoformer (Wu et al., 2021) extended Transformer with efficient attention and auto-correlation mechanisms respectively, while FedFormer (Zhou et al., 2022) and PyraFormer (Liu et al., 2022) utilized improved attention to achieve lower complexity. However, a recent study (Zeng et al., 2023) argued that Transformer may fail to correctly understand multivariate time series structure. PatchTST (Nie et al., 2022) address the issue with independent input but fell short in modeling multivariate correlations. CrossFormer (Zhang and Yan, 2023) tried to address with 2D attention, but its hierarchical segmentation failed adapt length per series. To address the challenges, we propose ARM to enhance the accuracy and adaptability of LTSF with proper training and data processing methods.
Figure 2: Overall Architecture of ARM with Vanilla Transformer as encoder-decoder. The left side depicts the global workflow and the right side illustrates the specific process in the model training.
@@ 3.1 Module A: Adaptive Univariate Effect Learning
In previous LTSF Transformers such as Informer and Autoformer (Wu et al., 2021; Zhou et al., 2021), the $L_{P}$ region of the decoder input is filled with 0 values as default outputs (models without an $L_{P}$ part input can also be considered to use 0 default). However, the actual meaning of 0 can vary significantly across series due to their distributional differences, making it difficult for multivariate models to determine the output level for each series. As the output level predominantly influences the loss like MSE during training, the model tends to use the majority of parameters to learn the output level, making it challenging to capture finer-grained temporal patterns and inter-series dependencies.
@@
Recent studies tried to mitigate the aforementioned issues by focusing on “eliminating distribution shifts between input and output”. These approaches attempt to estimate the mean and variance, eliminating them from the model input, and restoring these values at the output. RevIN (Kim et al., 2022) calculates the mean and variance for each input series while NLinear (Zeng et al., 2023) simply employs the last value of each series at $L_{I}$ as the output mean. However, as illustrated in Figure 1 (a), for series with differing statistical properties, the lookback lengths required to determine their output distribution tend not to be constant. Relying solely on last or all input values will weaken performance. Thus, we propose an adaptive method based on a learnable exponential moving average (EMA), which dynamically adjusts the output mean and variance for different series.
@@
Previous methods still employ constant value to fill the $L_{P}$ part of each series without building temporal patterns. In multivariate LTSF, handling the univariate temporal patterns prior to the multivariate modeling can help to learn inter-series relationships more accurately. In multivariate datasets, suppose a given series $i$ has no relationships with other series. If we utilize a univariate model to initialize $X_{P}^{i}$ part based on $X_{I}^{i}$ before other operations, this preliminary step will allow the model to focus more on capturing other inter-series relationships. Previous works like LSTNet (Lai et al., 2018) and ES-RNN (Smyl, 2020) have used classic autoregressive and exponential smoothing models for this purpose. However, both methods are limited in capturing basic patterns and suffer from poor generalization. Therefore, we use an adaptive method based on the Mixture of Experts (MoE) for temporal patterns to better handle these complexities, as illustrated in Figure 3 (c).
Figure 3: This figure illustrates six LTSF model types, categorized as univariate (Methods a-c) and multivariate (Methods d-f). The figure’s parallel arrows show independent series processing, and converging arrows indicate series mixing. Univariate models (a-c) process series separately: (a) employs individual models like DLinear for each series, (b) uses shared-parameter models such as PatchTST for all series, and (c) dynamically chooses best models per series, as in §3.1. Multivariate models (d-f) capture inter-series dependencies: (d) uses standard structures as Informer and Autoformer, outperformed by univariate methods; (e) combines univariate models with multivariate factors to better build inter-series dependencies but adding computational complexity; (f) introduces our Random Dropping strategy (see §3.2) for efficient learning of inter-series relationships.
@@
where $\circ$ denotes element-wise multiplication, $\bm{t}=[1,2,\cdots,L_{I}]$ is a vector representing the timesteps, and $\cdot]_{j}$ indicates the $j$-th element of the vector enclosed in brackets; $X_{I}^{(L_{I}-w_{j}:L_{I}),i}$ is a vector composed of the part of input series $X_{I}^{i}$ that goes back $w_{j}$ steps from the last value. When the $\alpha^{i}$ of channel $i$ approaches 1, $p_{global}^{i}=1$, and $p_{local}^{i}=0$, the AUEL of output distribution becomes RevIN. When $\alpha^{i}$ approaches 0, the AUEL will be similar to NLinear which uses last values.
AUEL of Temporal Patterns Figure 3 parts (a) and (b) illustrate two types of univariate input LTSF models: (b) with fully-shared parameters fits datasets with similar series characteristics while (a) with independent channels is suited data with diverse series. Typically, a dataset may contain multiple clusters of similar series. Thus, (c), an intermediary approach between (a) and (b) is preferable. We use MoE, simillar to the module in Switch Transformer (Fedus et al., 2022), to dynamically select the predictor based on series characteristics to initialize the temporal patterns for each series (see A.6.1 for more details on the implementation of MoE for temporal pattern learning).
Preprocessing and Inverse Processing After extracting $E^{i}$ and $S^{i}$ based on $X_{I}^{i}$, we perform the AUEL of output distribution upon the input series $X^{i}=\left[X_{I}^{i}\ \ \mathbf{0}_{p}^{i}\right]$, where $\mathbf{0}_{p}^{i}$ is the 0-filled $X_{P}$, and $\left[A\ \ B\right]$ denotes the concatenation of multivariate time series $A$ and $B$ along the time dimension, yielding the preprocessed $\widetilde{X}^{*i}$. Simillar to instance normalization (Ulyanov et al., 2016), we further apply a trainable channel affine transformation to each channel. Further, we utilize a MoE to obtain the final input $\widetilde{X}^{i}$. In this context, the MoE takes an input of length $L_{I}$ and produces an output of length $L_{P}$. We use $\gamma^{i}$, $\beta^{i}$ for trainable affine parameters, and $\epsilon$ for a small value to avoid overflow.
$\widetilde{X}^{*i}=\gamma^{i}\left(\frac{X^{i}-E^{i}}{S^{i}+\epsilon}\right)+\beta^{i},\ \ \ \ \widetilde{X}^{i}=\left[\widetilde{X}^{*i}_{I}\ \ \ \ \mathtt{MoE}\left(\left[\widetilde{X}^{*i}_{I}\ \ \ \ \widetilde{X}^{*i}_{P}\right]\right)\right]\vskip-5.0pt$ (2)
@@ 3.2 Module R: Random Dropping and Inter-Series Dependency Learning
In LTSF, accurately modeling inter-series dependencies is often challenging due to the simultaneous handling of complex temporal and contextual information. Figure 1 (c) provides an illustrative example where a multivariate Transformer fails to capture some very evident inter-series dependencies.
@@
We introduce a Random Dropping to train multivariate LTSF models effectively. This strategy facilitates the learning of inter-series dependencies by efficiently decoupling relationships between different series, without increasing computational complexity. Applied during the training phase, Random Dropping simultaneously sets a random subset of series to zero in both the input and training target, as illustrated in Figure 3 (f). The model thus learns the contributions of the current subset to the forecasting of series within that subset. Through this random selection, the model incrementally discerns which series significantly influence the forecasting of others, thereby effectively mitigating the risk of overfitting. Conceptually, Random Dropping can be viewed as a form of model ensemble that constructs a pool of forecasting models for every possible subset-to-subset relationship, then enabling the more effective models among them. Utilizing Random Dropping can straightforwardly enhance the learning ability of multivariate LTSF models in terms of inter-series dependency, achieving disentanglement between series.
@@ 3.3 Module M: Multi-kernel Local Smoothing
Figure 4: Structure of Multi-kernel Local Smoothing (MKLS). The central part of the figure illustrates the computation of MKLS, incorporating multiple 1D convolutions and channel attention. The left side and right side presents the application method of Pre-MKLS and Post-MKLS, respectively.
@@
In order to enhance the understanding of multivariate temporal structure, We propse the Multi-kernel Local Smoothing (MKLS) block which is used in conjunction with the Transformer blocks, as shown in Figure 4. MKLS uses multiple different 1D convolutional kernels and a channel-wise attention to learn and extract local information with adjustable view length. To address the aforementioned challenges of building reasonable temporal representations and enhancing locality, we provide two usage of MKLS. (i) Pre-MKLS is applied before feeding data into the Transformer blocks, Ahelping build meaningful local representations for multivariate input. (ii) Post-MKLS processes data in parallel with the Transformer blocks, playing a role like local attention with adjustable local windows. Through the incorporation of channel-wise attention, MKLS learns different kernel weights, providing robust adaptability for series with strong characteristic differences.
@@
MKLS block Let the input for the MKLS block be $X\in\mathbb{R}^{l\times d}$, where $l$ is the length of the temporal dimension and $d$ is the number of channels. Suppose we set $n$ 1D convolutional layers with kernel sizes $s_{j}\in\left(s_{1},\cdots,s_{n}\right)$. First, we obtain the computation results $\widetilde{X}_{j}$ from the $j$-th convolutional layer by: $\widetilde{X}_{j}=\mathtt{Conv1d}_{j}\left(X\right)$. To acquire the weights $m^{i}_{j}$ for kernel $j$ on channel $i$, we first perform kernel dropout and average-pooling on $\widetilde{X}_{j}$, and then calculate the cross attention between the averaged output and the original $X$. For kernel dropout, we randomly drop a certain proportion $r_{k}$ of kernel outputs, which enable independent learning of local patterns thus prevent overfitting. Let the tensor $\widetilde{\mathcal{X}}\in\mathbb{R}^{l\times d\times n}$ contain the results of each $\widetilde{X}_{k}\in\mathbb{R}^{l\times d}$. We now perform average-pooling to obtain another input $\widetilde{X}_{\mathtt{Avg}}=\frac{1}{1-r_{k}}\cdot\frac{\sum_{k=1}^{n}\widetilde{X}_{k}}{n}$ for the attention calculation besides original $X$. Then, we use $\widetilde{X}_{\mathtt{Avg}}$ and $X$ as Q and K in the attention, and obtain the kernel weights $M\in\mathbb{R}^{d\times n}$ for each channel using a trainable latent V. The specific calculation is: $M=\mathtt{softmax}\left(\frac{W_{Q}(X^{\top}+P_{E})\left(W_{K}(\widetilde{X}^{\top}_{\mathtt{Avg}}+P_{E})\right)^{\top}}{\sqrt{d}}\right)W_{V}$ , where $W_{Q}\in\mathbb{R}^{d\times d}$ and $W_{K}\in\mathbb{R}^{d\times d}$ are the weights for the Q and K in the attention, initialized as identity matrices to retain the original channel structure before training; $W_{V}\in\mathbb{R}^{d\times n}$ is designed to be a latent matrix, which projects the cross-channel attention map to the dimension of $n$, obtaining the kernel weights $M$; $P_{E}$ is the trainable positional embedding matrix. Now, we can calculate the weighted kernel output $\widetilde{S}$ as: $\widetilde{S}_{t,i}=\sum_{j=1}^{n}\left(\widetilde{\mathcal{X}}_{t,i,j}*M_{i,j}\right)_{t,i,j}+X_{t,i}$, where * represents the element-wise multiplication performed on the channel and multi-kernel dimensions; the output $\widetilde{S}_{t,i}$ is the local smoothed output after weighting $n$ local kernels. A residual shortcut of the input is added and denoted by $X_{t,i}$.
@@
Pre-MKLS and Post-MKLS MKLS module, used alongside Transformer blocks, effectively manages local shape extraction and alignment in multivariate input. Pre-MKLS constructs local smoothed representations, offering more self-adjustment based on input compared to the pure 1D convolutional approach like (Li et al., 2019). Post-MKLS, processing data in parallel with the Transformer, enables the latter to focus more on learning long-term connections, with its logic similar to the local attention methods such as (Child et al., 2019; Roy et al., 2021) and parallel local convolutional layer in (Wu et al., 2020), but with adjustable local view. The using methods of Pre-MKLS and Post-MKLS are shown on the left side and right side of Figure 4, respectively.